The physical fitness of an athlete is often measured by how much oxygen the athlete takes in (which is recorded in milliliters per kilogram, ml/kg). The mean maximum oxygen uptake for elite athletes has been found to be 70 with a standard deviation of 7.4. Assume that the distribution is approximately normal.

(a)  What is the probability that an elite athlete has a maximum oxygen uptake of at least 60 ml/kg?

answer:

(b)  What is the probability that an elite athlete has a maximum oxygen uptake of 70 ml/kg or lower?

answer:

(c)  Consider someone with a maximum oxygen uptake of 30 ml/kg. Is it likely that this person is an elite athlete? Write "YES" or "NO."

answer:
Answer #1

let x be amount of oxygen uptake for clite athlety. X~ N( 70, 74), u= 70,5=7.4. BY CLT, Ža ~ N (od X-70 7:4 a) ll x-70 7-4 41e) l( X5 30) ill x - 70 < 30-70 ༠) :Plzs - 5.41) 7.4 نو 19 1 - 0(5-41) = 1-1=0 f (x 30) = 0. So, no, the person is not an elifor further query please comment below.thank you